# Top Score on the Wrong Exam: On Benchmarking in Machine Learning for Vulnerability Detection

### v1

@@ 39 HTTP_Clone(&destination, source);
Figure 9: Context-dependent security: In this artificially crafted context, the function HTTP_Clone contains a null pointer dereference vulnerability. However, it is labeled as ’secure’ in the DiverseVul dataset because in its original context (see https://github.com/varnishcache/varnish-cache/commit/c5fd097e) it does not contain a security vulnerability.
Dependence on External Functions. The first and most prevalent type of dependence we identified is dependence on external functions. Consider the example in Figure Figure 8a. The heap-based buffer overflow in lines 20-21 of this function from the BigVul dataset depends on the external function cfg80211_find_vendor_ie. Without knowing this external function, we do not know what values vs_ie can have, and consequently, we do not know whether the buffer overflow can ever be triggered. In our empirical study, 43%-53% of context-dependent vulnerabilities can be attributed to dependence on external functions.
Dependence on Function Arguments. The second type of dependence we identified is dependence on function arguments. Consider the example in Figure Figure 8c. The null-pointer-dereference vulnerability in line 8 of this function from the BigVul dataset depends on the function argument sprinc. However, without knowing the context in which this function can be called, we do not know whether sprinc can ever be NULL. For example, it could be properly validated before passing it to the function in all cases where this function is actually called. In that case, there would be no vulnerability. In our empirical study, 39%-47% of context-dependent vulnerabilities can be attributed to dependence on function arguments.
Dependence on Type Declarations. The third type of dependence we identified is dependence on type declarations. Consider the example in Figure Figure 8b. The integer overflow vulnerability in line 20 depends on the type declaration of PP_Flash_MenuItem. Only if sizeof(PP_Flash_MenuItem) * menu->count exceeds the range of menu->items, the operation overflows. In our empirical study, 0-3% of context-dependent vulnerabilities can be attributed to dependence on type declarations.
Dependence on Globals. The fourth type of dependence we identified is dependence on globals, such as macros and global variables. Consider the example in Figure Figure 8d. Line 7 of this function from the Devign dataset is the cause of multiple buffer overflows in the JasPer repository (CVE-2014-8158). If HAVE_VLA is not defined and QMFB_JOINBUFSIZE is smaller than bufsize but still used without successful dynamic allocation, any attempt to use joinbuf for storing data will result in writing beyond its allocated size (QMFB_JOINBUFSIZE). In our empirical study, 2-10% of context-dependent vulnerabilities can be attributed to dependence on globals.
RQ.1-c Context-dependent Security
Inspired by our finding that the vulnerability of all 152 vulnerable functions in our empirical study could not be detected without considering additional context, we were also interested whether the same is true for secure functions. In other words, what proportion of non-vulnerable functions could be vulnerable if an appropriate external context existed?
Methodology. From each of the three datasets (BigVul, Devign, and DiverseVul), we randomly selected 30 samples that were labeled as secure. We publish the reproducible script and the sampled functions as part of our artifact. To find out what proportion of these functions could be vulnerable if an appropriate external context would exist, we tried to construct an artificial vulnerable context for each of them. Consider the example in Figure 9. In this artificially crafted context, the function HTTP_Clone (marked in green) contains a null pointer dereference vulnerability. However, it is labeled as ’secure’ in the DiverseVul dataset because in its original context [^1] https://github.com/varnishcache/varnish-cache/commit/c5fd097e it does not contain a security vulnerability. Without knowing the context, we can not decide whether this function is secure. Similar to this example, we manually tried to construct a vulnerable context for all 90 functions in our sample. For full transparency, we include the generated vulnerable settings (as explanations in natural language) in our artifact.
RQ.1 Result Summary
Within our sample of functions, about half of those labeled as vulnerable did not actually contain security vulnerabilities. All of those that do are context-dependent vulnerabilities, i.e., these functions are vulnerable only because an appropriate context exists under which the function is vulnerable. Similarly, the majority of secure functions contain context-dependent vulnerabilities that would make that function vulnerable if an appropriate context existed. As there is evidently insufficient information in the base units—that are used for training, validation, and testing in these datasets—to decide their vulnerability without further context, we conclude that ML4VD cannot be soundly evaluated as a classic binary classification problem on the function-level.
Problem statement (RQ.1). ML4VD techniques cannot be soundly evaluated as a classic binary classification problem on base units, such as functions.
@@
Since it is impossible for most functions to decide without further context whether they contain a vulnerability, there is currently no evidence that ML4VD techniques are actually capable of identifying security vulnerabilities in functions. But why do ML4VD papers still report high accuracy when evaluating their techniques using function-level datasets? What happens if we hide the code and only expose some features of the code, such as word counts?
Methodology. According to the CodeXGLUE benchmark for vulnerability detection [43], the state-of-the-art reported accuracy for the Devign dataset is 69%, achieved by the UniXcoder technique. UniXcoder [44] is a publicly available [^2] https://github.com/microsoft/CodeBERT/tree/master/UniXcoder large language model (LLM), pre-trained on 2.3 million functions and finetuned on the training subset of the Devign dataset for the task of vulnerability detection. During our initial experiments, we were able to reproduce the 69% accuracy using the evaluation subset of the Devign dataset. In order to test whether the Devign dataset contains spuriously correlated features that can be exploited to achieve a high accuracy without actually detecting security vulnerabilities, we then trained a simple classifier (Gradient Boosting Classifier) to detect vulnerabilities in Devign based on word counts only, completely disregarding the structure of the code. After training, we evaluated the resulting model on the evaluation subset of Devign to compute accuracy when no information on code structure and/or semantics is available during training.
@@
Results. The Gradient Boosting Classifier achieved an accuracy of 63.2% on the evaluation subset of Devign, which is only 5.8% lower than the state-of-the-art accuracy of UniXcoder (and within the Top-10 on CodeXGLUE). Surprisingly, for an effective vulnerability detection model, the whole process of training and evaluation only took 10 minutes on a MacBook Pro, which is extremely fast compared to state-of-the-art training times for LLMs. State-of-the-art LLMs require expensive Hardware (GPUs), weeks of computing time to pre-train, and at least multiple hours of computing time to finetune. Additionally, the Gradient Boosting Classifier was only trained on the training subset of Devign (21k functions) without any pre-training. For comparison, the UniXcoder technique utilized 2.3 million functions during pre-training. These results show that it is possible to achieve high accuracy on a function-level dataset while completely disregarding the structure and semantics of the code.
@@
Spurious Features (RQ2). Using word counts only, we were able to achieve 63.2% accuracy on the Devign dataset with a simple Gradient Boosting Classifier. This result shows that function-level datasets can be exploited to achieve high accuracy without actually detecting security vulnerabilities.
@@
Our results provide an alternative explanation for the results of ML4VD techniques that were reported in the literature. While we did not prove that state-of-the-art ML4VD techniques actually achieve their high scores by relying on spuriously correlated features, we have shown that it is possible to exploit these datasets to achieve high accuracy without actually detecting security vulnerabilities.
@@ VI-A Internal validity
Selection Bias. The random sampling of 100 functions from each dataset (BigVul, Devign, and DiverseVul) for our empirical study could introduce selection bias. Although random sampling aims to create a representative subset, it is possible that our sample may not fully capture the diversity and characteristics of the entire dataset. To mitigate this, we ensured that our sampling method was strictly random and publish the reproducible script and the sampled functions as part of our artifact [^3] Artifact: https://github.com/niklasrisse/TopScoreWrongExam.
Manual Labeling Errors. Labeling functions as vulnerable or secure and distinguishing between context-dependent and context-independent vulnerabilities involves subjective judgment. To address this, we employed a cross-labeling process where 200 out of the 300 functions were independently labeled for vulnerability by a second software engineering researcher, achieving a Cohen Kappa value of 0.7, indicating substantial agreement. Discrepancies were resolved through discussion, but some human error may still be present. For full transparency, we publish all labels and explanations as part of our artifact.
@@ VI-B External validity
Generalizability. Our study focuses on three specific datasets (BigVul, Devign, and DiverseVul), which are widely used in the ML4VD community. However, the findings may not generalize to other datasets or real-world software systems. Future studies should replicate our methodology on additional datasets and real-world codebases to validate our conclusions.
Dataset Composition. The datasets analyzed primarily contain C code. Our findings might not generalize to other programming languages with different syntactic and semantic properties. Future research should include datasets from various programming languages to evaluate the broader applicability of our results.
@@ VI-C Construct Validity
Performance Metrics. We used accuracy as the primary metric for evaluating classifier performance. While accuracy is standard, it does not differentiate between false positives and false negatives, which can have different implications in security contexts. Future studies should consider additional metrics such as precision, recall, and F1-score to provide a more nuanced evaluation of ML4VD techniques.
Spurious Correlations. Our results suggest that ML4VD techniques might achieve high accuracy by exploiting spurious correlations rather than genuinely detecting vulnerabilities. While we demonstrated this using word counts, further research should investigate other spuriously correlated features that might be exploited.
@@ VII Discussion and Future Work
The goal of empirical evaluation in ML4VD is to assess the capabilities of specific techniques by putting them to the test on real-world data. However, if the evaluation is based on incorrect assumptions, the results become meaningless. Our findings demonstrate that this is precisely the issue with ML4VD. Techniques can achieve ‘perfect scores’ by exploiting spurious features, even without possessing the capabilities the test aims to measure. In our specific case, the only plausible explanation for the good performance of ML4VD techniques are spurious correlations.
@@
Implications for general ML. Are we truly measuring the effectiveness of ML techniques in solving the specific tasks we expect them to solve? The issue of spurious correlations extends beyond ML4VD to other areas of machine learning, where the influence of spurious correlations may be less obvious. Therefore, it is crucial to develop methods to identify and measure these correlations during the benchmarking of ML techniques to ensure that we are genuinely evaluating their performance in solving the tasks they are intended to address.
Implications for Program Analysis. The context-dependency problem is a strong case for interprocedural analyses over intraprocedural analysis. While the latter is a common problem statement in static analysis, it is evidently insufficient for tasks requiring a comprehensive understanding of code contexts. Interprocedural analyses, which consider interactions between multiple procedures, are better suited to address the context-dependent nature of security vulnerabilities.
@@
Abstention. What could be possible ways forward for ML4VD? A simple possible solution to the context-dependency problem of function-level vulnerability detection could be to cast it as binary classification with abstention. Given a function, an ML4VD technique could either decide the vulnerability of the function or abstain from this decision. However, the obvious disadvantage of this approach is that only context-independent security vulnerabilities could actually be detected, which appear to be very rare. The context-dependency labels generated by our empirical study could be used as a starting point to explore this direction.
Alternative Granularities. Another alternative approach to function-level vulnerability detection could be to rely on higher granularities, such as file-level, module-level, or even repository-level. However, higher granularities could suffer from the same problems as the function-level approach. For example, deciding whether a file, a module, or even a complete repository contains a security vulnerability might still depend on external context. Additionally, the classifiers might still rely on spurious correlations within the larger context, leading to misleading results. These potential issues need to be empirically verified in future work to determine if higher granularities offer a true advantage or merely shift the problems to a different level of abstraction.
@@
Context-conditional Classification. Due to the hierarchical nature of software, determining the vulnerability of a given base unit (e.g., function) may require considering all necessary contexts, such as the complete state of a repository at a given commit ID and all other dependencies that might exist outside of the code (e.g. system of the user, dependencies of the repository, etc.). In the past, this approach was practically infeasible due to limitations of existing approaches (e.g., context size of LLMs). However, recent advances might make this approach feasible in the near future. Since Devign, BigVul, and DiverseVul all include patch commit IDs, which allow for the reconstruction of complete context for a given function, they could be used to evaluate ML4VD based on this alternative problem statement. Future research is needed to find optimal ways to include this context into state-of-the-art techniques and to optimally make decisions based on large contexts.
@@
However, adding repository-level context alone may not resolve the benchmarking problem, as an ML technique that disregards this context could still appear to perform well. Addressing this challenge will require ensuring that the evaluation metrics and benchmarks actually capture the vulnerability detection capabilities of the techniques that are tested. The context-dependency explanations generated by our empirical study could be used as a starting point to explore this direction.
@@
Overcoming Classification. Even with higher granularities, the essence of ML4VD would still be binary classification: Given an input, decide whether the input is vulnerable or secure. An alternative to classification could be generation of context. A context-generating ML4VD technique could generate the conditions under which a context-dependent vulnerability would be an actual vulnerability. For example, given a function, ML4VD techniques may generate a complete executable program in which the function contains a security vulnerability. This generated context could then be compared with the actual context to see whether the vulnerability exists in the real context.

### v2

@@ 39 HTTP_Clone(&destination, source);
Figure 11: Context-dependent security: In this artificially crafted context, the function HTTP_Clone contains a null pointer dereference vulnerability. However, it is labeled as ’secure’ in the DiverseVul dataset because in its original context (see https://github.com/varnishcache/varnish-cache/commit/c5fd097e) it is not known to contain a security vulnerability.
@@
Since it is impossible for most functions to decide without further context whether they contain a vulnerability, there is currently no evidence that ML4VD techniques are actually capable of identifying security vulnerabilities in functions. But why do ML4VD papers still report high scores when evaluating their techniques using function-level datasets? What happens if we hide the code and only expose some features of the code, such as word counts?
@@
Results. Table III shows the results for RQ.2. The Gradient Boosting Classifier achieved a f1-score of 86% on the evaluation subset of BigVul, a f1-score of 62.2% and an accuracy of 63.2% on the evaluation subset of Devign, and a f1-score of 11.1% on the evaluation subset of DiverseVul. For Devign and BigVul, the observed performances are only 4.4% (Devign, accuracy), 5.7% (Devign, f1-score), and 10% (BigVul) lower than the performances of recently published state-of-the-art techniques. Surprisingly, for an effective vulnerability detection model, the whole process of training and evaluation only took 30 minutes on a MacBook Pro, which is extremely fast compared to training times of state-of-the-art ML-based techniques. State-of-the-art ML-based techniques require expensive Hardware (GPUs), weeks of computing time to pre-train, and at least multiple hours of computing time to finetune. Additionally, the Gradient Boosting Classifier was only trained on the training subset of the two datasets without any pre-training. These results show that it is possible to achieve high performance on the most popular function-level datasets while completely disregarding the structure and semantics of the code.
TABLE III: Spurious Correlations. The performance of a simple model (Gradient Boosting Classifier) trained on word counts only (no information on code structure and/or semantics available during training).
Dataset State-of-the-art Without Structure
BigVul (f1-score) 0.96 0.86
Devign (f1-score) 0.68 0.62
DiverseVul (f1-score) 0.48 0.11
Devign (accuracy) 0.68 0.63
For the DiverseVul dataset, the Gradient Boosting Classifier was not able to achieve a comparable performance to the state-of-the-art techniques. This suggests that the dataset may contain fewer exploitable spurious features. However, this is not definitive proof. What the result does show is that state-of-the-art performance cannot be reached using a Gradient Boosting Classifier based solely on word count features.
@@
Spurious Features (RQ2). Using word counts only, we were able to achieve 62.2% f1-score on the Devign dataset and 86% f1-score on the BigVul dataset with a simple Gradient Boosting Classifier. These results show that the top-most widely-used function-level datasets can be exploited to achieve high scores without actually detecting security vulnerabilities.
@@
Our results provide an alternative explanation for the results reported in the literature. While we did not prove that state-of-the-art ML4VD techniques actually achieve their high scores by relying on spuriously correlated features, we have shown that it is possible to exploit these datasets to achieve high scores without actually detecting security vulnerabilities.
@@ VI-A Internal validity
Selection Bias. The random sampling of 100 functions from each dataset (BigVul, Devign, and DiverseVul) for our empirical study could introduce selection bias. Although random sampling aims to create a representative subset, it is possible that our sample may not fully capture the diversity and characteristics of the entire dataset. To mitigate this, we ensured that our sampling method was strictly random and publish the reproducible script and the sampled functions as part of our GitHub repository: https://github.com/niklasrisse/TopScoreWrongExam.
Manual Labeling Errors. Labeling functions as vulnerable or secure and distinguishing between context-dependent and context-independent vulnerabilities involves subjective judgment. To address this, we employed a cross-labeling process where all 300 functions were independently labeled by two Software Security researchers, achieving substantial agreement (82%) for labeling functions as vulnerable or secure and almost perfect agreement (98%) for distinguishing between context-dependent and context-independent vulnerabilities. Discrepancies were resolved through discussion, but some human error may still be present. For full transparency, we publish all labels and explanations as part of our artifact.
@@ VI-B External validity
Datasets. Our study focuses on three specific datasets (BigVul, Devign, and DiverseVul), which are widely used in the ML4VD community (cf. Fig. 3). However, the findings may not generalize to other datasets or real-world software systems. Future studies should replicate our methodology on additional datasets and real-world codebases to validate our conclusions.
Programming languages. The datasets analyzed primarily contain C code. Our findings might not generalize to other programming languages with different syntactic and semantic properties. Future research should include datasets from various programming languages to evaluate the broader applicability of our results.
Problem Statements. Our study focuses on the most prevalent problem statement of ML4VD as function-level binary classification problem (cf. §III and Fig. 2) where the classifier makes the decision exclusively based on information within the given function. However, there exist other definitions of ML4VD (e.g., at other granularities, such as line- or file-level, and even some that explicitly consider context, such as vulnerability-specific inter-procedural slicing [38]). While we discuss our perspective for other definitions of ML4VD in Section VII, future studies should carry out similar investigations to validate how reasonable these problem statements are indeed.
@@ VI-C Construct Validity
Spurious Correlations. Our results suggest that ML4VD techniques might achieve high scores by exploiting spurious correlations rather than genuinely detecting vulnerabilities. While we demonstrated this using word counts, further research should investigate other features that might be spuriously correlated with the vulnerable label. We discuss some related work in Section II.
@@ VII Discussion and Future Work
The goal of empirical evaluation in ML4VD is to assess the capabilities of specific techniques by putting them to the test on real-world data. However, if the evaluation is based on incorrect assumptions, the results become meaningless. Our findings demonstrate that this is precisely the issue with ML4VD. Techniques can achieve ‘top scores’ by exploiting spurious features, even without possessing the capabilities the test aims to measure. In our specific case, the only plausible explanation for the good performance of ML4VD techniques are spurious correlations.
@@
Implications for general ML. Are we truly measuring the effectiveness of ML techniques in solving the specific tasks we expect them to solve? The issue of spurious correlations is not limited to ML4VD but extends to other areas of machine learning, where it has been explored in depth by several studies [39, 40]. It is crucial to continue the development of methods to identify and measure these correlations during the benchmarking of ML techniques to ensure that we are genuinely evaluating their performance in solving the tasks they are intended to address.
Implications for Program Analysis. The prevalence of the context-dependency problem suggests that an effective function-level program analysis, such as an intra-procedural static analysis, must consider a reasonable approximation of the space of valid external states during the average function call, e.g., encoded as a function precondition, function summary, or type system. While the context-dependency check can ultimately be delegated to the user in practice, given our results, we suggest that particular care is taken when evaluating such tools at the function-level in the absence of users. In addition to function-level program analysis, there are inter-procedural, or whole-program analyses, which consider interactions between multiple procedures, and thus implicitly resolve the context-dependency problem during the search for security vulnerabilities.
@@
Abstention. What could be possible ways forward for ML4VD? A simple possible solution to the context-dependency problem of function-level vulnerability detection could be to cast it as binary classification with abstention. Given a function, an ML4VD technique could either decide the vulnerability of the function or abstain from this decision. However, the obvious disadvantage of this approach is that only context-independent security vulnerabilities could actually be detected, which appear to be very rare (cf. RQ.1-b). The context-dependency labels generated by our empirical study could be used as a starting point to explore this direction.
Other Base Units. Instead of deciding whether a function is vulnerable, we could decide whether other types of base units are vulnerable, such as lines, statements, files, modules, commits, or vulnerability-specific inter-procedural slices (cf. Fig. 2). However, for smaller base units, such as line-, statement-, or commit-level, there is no reason to believe that the context-dependency problem is resolved (subject to further study) while for larger base units, such as files or modules, the context-dependency problem may be much less pronounced (subject to further study), although the vulnerability label at that such coarse granularity may be less actionable for the security researcher. For alternative granularities, future work shall determine if they offer a true advantage or merely shift the problem to a different level of abstraction.
The vulnerability-specific inter-procedural slice [38] (at least conceptually) includes all required context to trigger a vulnerability of a specific type at a specific code location, and thus resolves the context-dependency problem. However, it also somewhat delegates the detection problem because it requires, as slicing criterion, a (vulnerability-specific) statement-of-interest where a vulnerability may be triggered. In general, we believe that a combination of static analysis and machine learning [41] offers a promising avenue for future research in vulnerability detection.
@@
Context-conditional Classification. We could define ML4VD as a base-unit-level classification problem given, e.g., the entire software repository as context. Due to the hierarchical nature of software, determining the vulnerability of a given base unit (e.g., function) may require considering a large scope of context, such as the complete state of a repository at a given commit ID and all other dependencies that might exist outside of the code (e.g. system of the user, dependencies of the repository, etc.). Since the Devign, BigVul, and DiverseVul benchmarks all include patch commit IDs, the complete context for a given function could be reconstructed, so that they could be used to evaluate ML4VD based on this alternative problem statement. Future research is needed to find optimal ways to include this context into state-of-the-art techniques.
@@
However, adding repository-level context alone may not resolve the benchmarking problem. Evidently, ML4VD techniques that disregard this context still appear to perform well (likely due to spurious correlations; RQ.2). Addressing this challenge will require ensuring that the evaluation metrics and benchmarks actually capture the vulnerability detection capabilities of the techniques that are tested.
@@
Overcoming Classification. Even using different types of base units and providing the entire repository as context, ML4VD would still be a binary classification task: Given a unit, decide whether the unit is vulnerable or secure. Alternatively, an ML4VD technique could provide the condition under which a given function would actually be vulnerable, e.g., by generating an example context: Given a function, an ML4VD technique may generate a complete executable program which is vulnerable to attack due to the code in that function. This generated context could then be compared with the actual context to see whether the vulnerability exists in the real context. A much stronger approach could be the (ML-assisted) generation of a proof of vulnerability in form of a test case, which demonstrates the exploitation of the vulnerability in the given context.
@@
VIII Data Availability
To ensure the reproducibility of our results and to provide transparency in our research, we have made all related scripts and data publicly available. All resources can be accessed as part of our GitHub repository, which is available at https://github.com/niklasrisse/TopScoreWrongExam.
Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972. This research is also partially funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is supported by ERC grant (Project AT_SCALE, 101179366).